OpenStudy (anonymous):

solve 2sin^2 x - sin x = 1 for 0 less than or equal to c less than 360 degrees

OpenStudy (cwrw238):

2 sin^x - sin x - 1 = 0 (2sin x + 1)(sin x - 1) = 0 so sin x = -1/2 , 1

OpenStudy (anonymous):

so only two solutions?

OpenStudy (cwrw238):

only 2 solutions for sin x but u want values of angle x between 0 and 360 right?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

angle whose sine = 1/2 is 30 degrees but its -1/2 so its in 3rd and 4th quadrants so x = 210 , 330 degrees

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so 7pi/6 and 11pi/6 right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

okay :)

OpenStudy (cwrw238):

arcsin 1 = 90 degrees or pi/2 so complete solution is x = pi/2 . 7pi/6 and 11pi/6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

will i do this the same way you showed me? tan^2 x - 3 = 0 for 0 less than or equal to x less than 2pi

OpenStudy (cwrw238):

yes find values of tan x, remember square root can be positive or negative. and the tangent is positive in first and third quadrants and negative in the others

OpenStudy (anonymous):

ok